A Bibliography of Publications about the Data Explorer Scientific Visualization System

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org, beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

and

Keith Sams

IBM Visualization Data Explorer 1507 LBJ Freeway, Dallas, TX 75234, USA

e-mail: ksams@ net.ibm.com

12 August 2013 Version 0.24

Title word cross-reference

10th [Dat94]. 11th [Ano95a]. 160th [Hea94].

2 [Ano94a]. 2-D [Ano94a].

3 [Ano94a]. 3-D [Ano94a]. 3D [ACF^{+}94, BMN93].

4d2 [SN95]. 4th [SC95].

6th [Hal94].

90’s [Bur93]. ’91 [NR91]. ’92

[KN92b, KN92a]. ’93 [Sup93, NB93]. ’94 [Hal94, Hea94, OS94]. ’95 [NS95, Sca95].

AAAS [Hea94]. Access [CCT93, TKFG93]. Adding [Watxx]. Advanced [N^{+}94]. Advancement [Hea94]. Advances [IEE96, Phi96]. Aeolian [Pic94]. Agricultural [F^{+}94]. Aided [IEE96, PMA^{+}95, Phi96]. Alpha [Ano94a]. American [Hea94]. Analysis [AT95a, AT95c, Kna93, Kna94, Tre92a, Tre93h, Tre93i, Tre95a, AT95b, ACC94, Tre93c]. ANDECS [FG96]. Animation [EW93, GW95, KCWG95]. Annual [Ano95b, Saf94]. Anthropology [K^{+}92]. application [LNH^{+}95]. Applications [AFMS96, Hos94, Saf94, EW93]. applying [HM95a]. Approach [Kna94, Tre95a]. April [LCHS96]. Architecture [AT95a, AT95c, L^{+}92, RT93b, RT93a, AT95b, LAC^{+}92, RT93d]. architectures [Hal94]. Archives [Tre93d, Tre94a, Tre94b, Tre95b, Tre94e]. Art [Phi96]. Artificial [Hos94]. aspects [N^{+}94]. Assisting [BRT95]. Association [Hea94]. Athens [Hal94]. Atlanta [NS95]. Atmospheric [RBP^{+}92, Tre94h, Tre95a]. Automatic [Pic94].

Based [F^{+}94, RT93b, RT93a, BRT95, BMN93, RT93d]. Belgium [LCHS96]. between [FTD95, HPM^{+}00]. Biochemical [KG96]. Blackboard [HM95b, TC96]. Block [Lan94]. Boston [KN92b, KN92a, Sch93]. British [Ano95b]. Browsing [Tre93d]. Brussels [LCHS96].

CA [FSW94, NR91, NB93]. CACSD [FG96]. Cagliari [Sca95]. calculations [PMA^{+}95]. California [RA94, SC95]. Canada [Ano95b]. Care [Saf94]. Case [HM95a, RBP^{+}92]. Center [Hos94]. Centered [Saf94]. CFD [SN95]. Changing [Hea94]. Chemical [LG93]. Climatology [Tre92b, Tre93e]. CO [Bur93]. Coastal [KCWG95, SC95]. Cognitive [N^{+}94]. Collins [Bur93]. Colloquium [IEE96]. Colormap [BRT95]. Columbia [Ano95b]. communication [HPM^{+}00]. Communiqué [Com93]. Composite [D^{+}93]. Computer [IEE96, Lan94, PMA^{+}95, Saf94, N^{+}94, Phi96]. Computer-Aided [IEE96, PMA^{+}95, Phi96]. Computing [Ano94b, GW95, SHP94, Saf94, FSW94, LCHS96]. Conference [Ano95a, FSW94, Hal94, Hos94, Dat94, SHP94, LCHS96, SC95, Ano94b]. conjunction [Phi96]. Consumer [ACGP93]. contaminant [MJS98]. Control [F^{+}94, IEE96]. Correlative [TG91, Tre92a, Tre93f, Tre93c]. Cover [SB94]. Cranial [D^{+}93]. CT [BMN93, D^{+}93].

D [Ano94a]. Dallas [Ano95a]. Data [AT95a, AT95c, BO95, ECSK95, FTD95, GW95, H^{+}91, HMM94, HMR95, Com93, Dat94, Kna94, RT93c, RT98, Tre92a, Tre92c, Tre93d, TKFG93, Tre93g, Tre93a, Tre93h, Tre93i, Tre94a, Tre94b, Tre94c, Tre94f, Tre94i, Tre94d, Tre95a, Tre95b, Tre95c, Watxx, AT95b, Hau93, MJS98, SN95, TG91, Tre93c, Tre94e, Tre94g, ACC94, Ano94a, B^{+}xx, GR95, IBM93a, IBM93b, Lan94]. Data-Flow [AT95a, AT95c, AT95b]. Data-Parallel [HMM94]. Dataflow [Lan94, R^{+}92, RBM^{+}92, DL96]. Datasets [CCT93, LGH96]. DBMS [ACF^{+}94]. DC [OS94]. Debugging [AFMS96]. defined [Watxx]. Demands [Bur93]. Depletion [Tre93b]. Design [F^{+}94, IEE96]. Development [HM94]. Diagram [Lan94]. Diego [NR91, SC95]. Digest [IEE96]. digital [RA94]. Dimensional [BW93, D^{+}93]. Disparate [Tre94i]. Display [ECSK95, HMR95, Tre94f, RA94]. Distribution [Tre94f, LGH96]. domestica [HPM^{+}00]. dramatic [Phi96]. Dynamics [KCWG95, Tre94h, FG96].

Early [KDHB94]. Earth [Tre92a, Tre93f, Tre94i, Tre95a, Tre95b, Tre93c]. Ecological [FRN94]. Education [Ano94b, FSW94]. Electronic [AR93, PMA^{+}95]. Electrostatics [GR95]. Element [Watxx]. end [RT98]. Engineering [Dat94].

Environment [HM94, LNH^{+}95]. Environmental [RBP^{+}92, Tre93h, Tre95a]. environments [FTD95]. Enzyme [GR95]. estuaries [MJS98]. Estuarine [SC95]. Europe [Hal94, LCHS96]. Evaluation [F^{+}94]. Events [Tre94c, Tre94d, Tre94g]. Examination [Tre94f]. exhibition [LCHS96]. Expert [F^{+}94, ACC94]. Explorer [Com93, ACC94, Ano94a, B^{+}xx, GR95, IBM93a, IBM93b, Lan94, Wat94, Watxx]. Extended [AT95a, AT95c, AT95b]. Extending [ACF^{+}94].

February [Hea94, Dat94, KL93, RA94]. Federal [Bur93]. Fingering [BW93]. Flight [Hos94]. Flow [AT95a, AT95c, GW95, KCWG95, AT95b]. forecasting [TC96]. Fort [Bur93]. Fossils [K^{+}92, KDHB94, KDH95]. fractals [Pic94]. framework [FG96]. Francisco [Hea94]. Frontiers [KG96]. Fusion [BMN93, Sch93].

GA [NS95]. ganglionaris [HPM^{+}00]. Generalized [Tre95a]. Generation [HM94, Pic94]. Geographic [Ano95b, Kna93, Kna94, N^{+}94]. GIS [FTD95, F^{+}94]. Global [Tre92b, Tre93e, Tre93i, Tre94f]. Goddard [Hos94]. Graphics [Lan94, PMA^{+}95]. Greece [Hal94]. Greenbelt [Hos94]. Grids [H^{+}91]. Guide [ACGP93].

held [Hos94]. High [SHP94, LCHS96, SN95]. High-Performance [SHP94, LCHS96, SN95]. Homo [KDHB94]. Houston [Dat94]. HPCN [LCHS96]. Human [K^{+}92, RA94, KDH95, N^{+}94]. human-computer [N^{+}94]. Hydrologic [Bur93, HK93]. Hydrology [Ano95a].

i.e [Hos94]. IBM [Ano94a, IBM93a, IBM93b, Pic94, Watxx]. IFIP [Ano94b, FSW94]. Illuminate [Tre94i]. Images [ACF^{+}94, BMN93]. Imaging [AR93]. Information [Ano95a, Ano95b, N^{+}94]. Innovation [BO95]. Insight [BO95]. Integrates [ECSK95]. integrating [FG96]. Intelligence [Hos94]. interaction [N^{+}94]. Interactive [Ano95a, RT94, Tre94e, Tre94a, Tre94b, Tre95b, FG96]. Interagency [Bur93]. interchange [FTD95]. International [Ano95a, Hal94, Dat94, Sca95, LCHS96, SC95]. interpretation [PMA^{+}95]. Intrepid [Wat94]. Irregular [H^{+}91]. Irvine [FSW94]. Italy [Sca95].

January [Ano95a]. Jersey [MJS98]. Joint [KL93]. Jose [NB93, RA94]. July [Hal94]. June [Bur93].

Knowledge [F^{+}94]. Knoxville [SHP94].

lamina [HPM^{+}00]. Landmark [BMN93]. Landmark-based [BMN93]. Language [Lan94, DL96]. languages [Hal94]. Large [CCT93, Tre93d, Tre94f]. libraries [LGH96]. Lidar [ECSK95]. Lie [RT96]. loops [FG96].

MA [KN92a, Sch93]. Mallorca [N^{+}94].

Management

[Ano95b, Tre93d, Tre93a, Tre94f, Tre92c]. Maps [ACGP93]. March [Ano95b, N^{+}94]. Maryland [Hos94]. Massachusetts [KN92b]. May [Hos94, SHP94]. Medical [ACF^{+}94, Phi96, Saf94]. Meeting [Hea94, KL93]. Meteorological [Tre94c, Tre94d, Tre95c, Tre94g]. Meteorology [Ano95a]. Methodology [AFMS96, B^{+}xx]. methods [Tre92c]. Miscible [BW93]. Model [GW95, H^{+}91]. Modeling [BW93, Bur93, FRN94, HK93, FG96, MJS98, SC95]. Models [TKFG93]. Modular [FRN94]. Molecular [BO95, PMA^{+}95]. Morocco [KDHB94]. Multidimensional [Tre93g, TG91]. Multiple [K^{+}92]. Musca [HPM^{+}00]. MVE [LNH^{+}95].

NASA [Hos94]. National [Hea94]. NATO [N^{+}94]. Natural [Ano95b]. networking [LCHS96]. neurons [HPM^{+}00]. Newsletter [Com93]. Next [Ano95b, HM94]. Next-Generation [HM94]. Ninth [Ano95b]. non [Watxx]. non-defined [Watxx]. Northwestern [Tre94c, Tre94d, Tre94g]. November [Sup93, NS95].

Object [R^{+}92, RBM^{+}92]. Object-Oriented [R^{+}92, RBM^{+}92]. Oceanography [Ano95a, Hau93]. October [KN92b, KN92a, NR91, NB93, NS95, OS94, SC95]. Olympics [TC96]. Oregon [Sup93]. Orientation [ACGP93]. Oriented [R^{+}92, RBM^{+}92]. Orleans [KL93]. Ozone [Tre92b, Tre93e, Tre93b, Tre94f, Tre94h].

packages [Ano94a]. Palma [N^{+}94]. Panel [RBM^{+}92]. Paradigm [R^{+}92, RBM^{+}92]. Parallel [HM94, HMM94, HMR95, HM95a, HM95b, Hal94, Pic94]. PARLE [Hal94]. particle [SN95]. Patient [Saf94]. Patient-Centered [Saf94]. Perceptual [RT93b, RT93c, RT94]. Performance [HM94, HMM94, HMR95, SHP94, HM95a, HM95b, LCHS96, SN95]. Peru [Tre94c, Tre94g, Tre94d]. physical [Hau93]. Pieces [K^{+}92]. Polar [Tre93b]. polymers [PMA^{+}95]. Portland [Sup93]. ports [Ano94a]. Power [GPT92, Pic94]. PRACTICAL [RT96]. Practices [F^{+}94]. Principles [Tre93a]. Proceedings [AR93, Ano95b, Bur93, Hea94, Sup93, Dat94, SHP94, FSW94, Hal94, Hos94, KN92b, KN92a, LCHS96, NR91, NB93, NS95, OS94, Sca95, SC95]. Processing [Ano95a, LGH96, RA94]. Program [ECSK95]. Programs [HMM94, HM95b]. prototype [ACC94]. Prototyping [HM94]. Provisions [H^{+}91].

QBISM [ACF^{+}94]. Quality [F^{+}94]. Quarry [D^{+}93].

rainbow [RT98]. Rainfall [Tre94c, Tre94d, Tre94g]. Randomly [SB94]. RCWP [F^{+}94]. real [Hau93]. Reality [GW95, KDHB94, KG96]. Reconstruction [D^{+}93, K^{+}92, KDHB94, KDH95]. Regular [H^{+}91]. Relative [AFMS96]. Remains [D^{+}93]. Renderer [Luc92]. Requirements [RBP^{+}92]. Research [KG96, N^{+}94]. Resources [Ano95b]. right [LNH^{+}95]. Rotations [ACGP93]. Rule [BRT95, RT93b, RT93a, RT93d]. Rule-Based [RT93b, RT93a, BRT95, RT93d]. Rules [RT94].

Sale [D^{+}93]. San [Hea94, NR91, NB93, RA94, SC95]. Sapiens [KDHB94]. Scalable [FRN94, HMM94, SHP94]. scanning [Phi96]. Scattered [Tre94c, Tre94d, Tre95c, MJS98, Tre94g]. Science [BO95, Hea94, Phi96, Tre95b]. Sciences [RBP^{+}92, Tre92a, Tre93f, Tre94i, Tre95a, Tre93c]. Scientific [AFMS96, Ano94b, B^{+}xx, CCT93, H^{+}91, LG93, Lan94, Luc92, L^{+}92, MJS98, Sca95, TKFG93, Tre93a, Tre94a, Tre94b, Tre94f, DL96, EW93, FSW94, HM95a, LAC^{+}92, Tre92c, Tre94e]. sediment [MJS98]. Selection [BRT95, F^{+}94]. Sensor [Sch93]. September [Sca95, Sch93]. Sets [Tre94f, Tre94i]. Seventeenth [Saf94]. Severe [Tre94c, Tre94d, Tre94g]. Shift [R^{+}92, RBM^{+}92]. simulated [Hau93]. simulation [FG96]. Smart [CCT93]. Software [TKFG93]. Solutions [Tre94f]. Space [Hos94, Tre93f, Tre93c]. Spain [N^{+}94]. SPECT [BMN93]. SPIE [AR93]. SPIE/SPSE [AR93]. SPSE [AR93]. statistical [ACC94]. Step [Ano95b]. Stratospheric [Tre92b, Tre93e, Tre93b, Tre94h]. structure [PMA^{+}95]. Structures [RT93c, TKFG93]. Study [RBP^{+}92, HM95a, Tre94g]. Supercomputing [Sup93]. Support [ACF^{+}94, LGH96]. Symposium [AR93, Ano95b, Saf94, Sca95]. Synapse [SB94]. System [FRN94, F^{+}94, GPT92, IEE96, L^{+}92, ACC94, LAC^{+}92, Pic94]. Systems [Ano95a, Ano95b, LG93, R^{+}92, RBM^{+}92, N^{+}94].

Task [Kna93, Kna94]. Teaching [Lan94]. Techniques [ACGP93, HK93, KCWG95, Tre94f, TG91, Tre93c]. technologies [Phi96]. Texas [Ano95a]. Thomas [D^{+}93]. Three [BW93, D^{+}93]. Three-dimensional [D^{+}93]. TN [SHP94]. Tool [BRT95, SN95]. Tools [BO95, EW93, FG96, Phi96]. TX [Dat94]. Types [Watxx].

Unifying [Tre93a]. University [Ano94b, FSW94]. USA [Bur93, FSW94, Dat94, SHP94, KN92a, NR91, NB93, NS95, OS94, Sch93]. Uses [Ano94b, FSW94]. Using [DL96, Lan94, RT94, B^{+}xx].

V [RA94]. Vancouver [Ano95b]. VI [Sch93]. Vipar [LGH96]. Virtual [GW95, KDHB94, KG96]. Viscous [BW93]. vision [RA94]. Visual [HMR95, Tre94f, RA94]. Visualization [AT95a, AT95c, Ano94b, BW93, BO95, GPT92, GR95, GW95, H^{+}91, HM94, HMM94, HM95b, Hau93, HK93, K^{+}92, KN92b, KN92a, KG96, Kna93, Kna94, LG93, Lan94, Luc92, L^{+}92, NR91, NB93, NS95, OS94, Phi96, Pic94, RBP^{+}92, R^{+}92, RBM^{+}92, RT93b, RT93a, RT94, RT96, Sca95, Tre93d, Tre93f, TKFG93, Tre93a, Tre93h, Tre93i, Tre93b, Tre93c, Tre94c, Tre94g, Tre94d, Tre94h, Tre95a, Tre95c, TC96, AT95b, Ano94a, DL96, EW93, FTD95, FSW94, HM95a, IBM93a, IBM93b, LGH96, LNH^{+}95, LAC^{+}92, MJS98, RT93d, RT98, SN95, TG91, Tre92c, RT96]. Visualizations [Tre94i]. Visualizing [ACGP93, GR95, HM95b, HPM^{+}00, KCWG95]. Vortex [Tre93b].

Walking [SB94]. Washington [OS94]. Water [F^{+}94]. Watersheds [F^{+}94]. Weather [TC96]. Web [DL96]. WG [FSW94]. WG3.2 [Ano94b]. Wide [DL96]. Working [Ano94b, FSW94]. Workshop [Bur93, N^{+}94]. World [Hea94, DL96].

York [MJS98]. York/New [MJS98].

References

Aliferis:1994:DEP

[ACC94] C. Aliferis, E. Chao, and G. F. Cooper. Data Explorer: a prototype expert system for statistical analysis. In Safran [Saf94], pages 389–393. ISBN 0-07-001502-3. LCCN W3 SY5074K 1993.

Arya:1994:QED

[ACF^{+}94] M. Arya, W. Cody, C. Faloutsos, J. Richardson, and A. Toga. QBISM: Extending a DBMS to support 3D medical images. In Data Engineering ’94 [Dat94], pages 314–325 (or 314–327??). ISBN 0-8186-5402-3 (case), 0-8186-5400-7 (paperback), 0-8186-5401-5 (microfiche). LCCN QA76.9.D3 I5582 1994. IEEE Computer Society Press order number 5400-02. IEEE Catalog Number 94CH3383-7.

Alpern:1993:OMT

[ACGP93] B. Alpern, L. Carter, M. Grayson, and C. Pelkie. Orientation maps:

Techniques for visualizing rotations (A consumer’s guide). In Nielson and Bergeron [NB93], pages 183–188. ISBN 0-8186-3942-3 (case), 0-8186-3940-7 (paper), 0-8186-3941-5 (microfiche). ISSN 1070-2385. LCCN T385 .I192 1993.

Abramson:1996:RDN

[AFMS96] David Abramson, Ian Foster, John Michalakes, and Rok Socič. Relative debugging: A new methodology for debugging scientific applications. Communications of the ACM, 39(11): 69–77, November 1996. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.cit.gu.edu.
au/~davida/guard.html. Uses

IBM Data Explorer to illustrate the computational differences between two versions of the same program, using the Guard (Griffith University Relative Debugger) system to automatically carry out run-time monitoring and comparison of results of two implementations of the same program, possibly written in different programming languages, and possibly running on different computer architectures at different geographic locations.

Anonymous:1994:IPD

[Ano94a] Anonymous. IBM ports its Data Explorer 2-D and 3-D visualization packages to Alpha. Digital News and Review, 11(16):8–??, August 1994. ISSN 1065-7452.

Anonymous:1994:VSC

[Ano94b] Anonymous. Visualization in scientific computing: Uses in university education. IFIP WG3.2 Working Conference. IFIP Transactions. A. Computer Science and Technology, A-??(????):??, ???? 1994. CODEN ITATEC. ISSN 0926-5473.

Anonymous:1995:ICI

[Ano95a] Anonymous, editor. 11th International Conference on Interactive Information and Processing Systems for Meteorology, Oceanography, and Hydrology: January 15– 20, 1995, Dallas, Texas, International conference on interactive information and processing systems for meteorology oceanography and hydrology. American Meteorological Society, Boston, MA, USA, 1995.

Anonymous:1995:NSN

[Ano95b] Anonymous, editor. The Next Step: Ninth Annual Symposium on Geographic Information Systems in Natural Resources Management, March 27–30, 1995, Vancouver, British Columbia, Canada. Symposium Proceedings. GIS World, Fort Collins, CO, USA, 1995. ISBN 1-882610-15-6. Two volumes.

Allebach:1993:PSS

[AR93] J. P. Allebach and B. E. Rogowitz, editors. Proceedings of the SPIE/ SPSE Symposium on Electronic Imaging, number 1913 in Proceedings — SPIE, The International Society for Optical Engineering. Society of Photo-optical Instrumentation Engineers (SPIE), Bellingham, WA, USA, February 1993. ISBN 0-8194-1146-9. ISSN 0361-0748. LCCN ????

Abram:1995:EDAb

[AT95a] G. Abram and L. A. Treinish.

An extended data-flow architecture for data analysis and visualization. In Nielson and Silver [NS95], pages 263–270, 461. ISBN 0-7803-3127-3, 0-8186-7187-4, 0-7803-3128-1. ISSN 1070-2385. LCCN T385 .I198 1995. IEEE Computer Society Press order number PR07187. IEEE Catalog Number 95CB35835.

Abram:1995:EDAa

[AT95b] Greg Abram and Lloyd A. Treinish. An extended data-flow architecture for data analysis and visualization. Research report RC 20001 (88338), IBM T. J. Watson Research Center, Yorktown Heights, NY, USA, February 21, 1995. 15 pp.

Abram:1995:EDAc

[AT95c] Greg Abram and Lloyd A. Treinish. An extended data-flow architecture for data analysis and visualization. Computer Graphics, 29(2):17–21, May 1995. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Bird:19xx:SMU

[B^{+}xx] C. L. Bird et al. Scientific methodology using Data Explorer. Report 278, IBM UK Scientific Center, IBM Hursley, Winchester, UK, 19xx. ???? pp.

Brown:1993:LFS

[BMN93] L. G. Brown, G. Q. Maguire, Jr., and M. E. Noz. Landmark-based 3D fusion of SPECT and CT images. In Schenker [Sch93], pages 166–175 (or 166–174??). ISBN 0-8194-1324-0. ISSN 0361-0748. LCCN TS510.S63 v.2059. Sensor Fusion VI.

Bowie:1995:DVM

[BO95] Jack E. Bowie and Arthur J. Olson, editors. Data Visualization in Molecular Science: Tools for Insight and Innovation. Addison-Wesley, Reading, MA, USA, 1995. ISBN 0-201-58157-4. xviii + 285 pp. LCCN QD461 .D22 1995. Contents: Introduction — Application visualization system — Quantum chemistry visualization using the AVS Chemistry Viewer — Visualizing enzyme electrostatics with IBM Visualization Data Explorer — Interactive visualization of molecular scenarios with MOLCAD/SYBYL — Molecular surface comparison using Anaconda — Visualization approaches in quantum chemistry using UniChem — Visualizing electrostatic interactions with Insight II — Cerius2 molecular modeling software for industrial applications — The CAChe WorkSystem and molecular mechanics: coordination of DBT on MoS2 — Visualization and representation in chemical education using HyperChem — Visualization supplier reference.

Bergman:1995:RTA

[BRT95] L. D. Bergman, B. E. Rogowitz, and L. A. Treinish. A rulebased tool for assisting colormap selection. In Nielson and Silver [NS95], pages 118–125, 444. ISBN 0-7803-3127-3, 0-8186-7187-4, 0-7803-3128-1. ISSN 1070-2385. LCCN T385 .I198 1995. IEEE Computer Society Press order number PR07187. IEEE Catalog Number 95CB35835.

Burton:1993:PFI

[Bur93] James S. Burton, editor. Proceedings of the Federal Interagency Workshop on Hydrologic Modeling Demands for the 90’s: June 6–9, 1993, Fort Collins, CO, USA, number 93-4018 in Water-resources investigations report. U.S. Geological Survey, Reston, VA, USA, 1993. ISBN ???? LCCN GB701 .W39 no. 93-4018. PB Reports PB95-186615.

Barley:1993:MVT

[BW93] J. J. Barley and D. W. Williams.

Modeling and visualization of three dimensional miscible viscous fingering. In Earnshaw and Watson [EW93], pages 261–278. ISBN 0-12-227745-7. LCCN T385 .A68 1993. Proceedings of the British Computer Society Conference on Animation and Scientific Visualization, December 1992.

Christian:1993:SAL

[CCT93] C. Christian, R. Cromp, and L. Treinish. Smart access to large scientific datasets. In Supercomputing ’93 [Sup93], pages 372–375. ISBN 0-8186-4342-0, 0-8186-4340-4, 0-8186-4341-2 (microfiche). ISSN 1063-9535. LCCN QA76.5 .S894 1993. IEEE Computer Society Press order number 4340-02. IEEE catalog number 93CH3342-3. ACM order number 415932.

IBM:1993:DEC

[Com93] Data Explorer Communiqué Newsletter, 1(1), April 1993. IBM Corporation, Yorktown Heights, NY, USA.

Dean:1993:TCC

[D^{+}93] D. Dean et al. Three-dimensional CT composite reconstruction of sale and Thomas quarry cranial remains. In Proceedings of Symposium on Image Acquisition, Analysis and Archiving: New Research Directions in Physical Anthropology, 62nd Annual Meeting of the American Association of Physical Anthropologists, April, 1993, page ?? ????, ????, 1993. ISBN ???? LCCN ???? N. American Journal of Physical Anthropology, Supplement 16:79.

IEEE:1994:PIC

[Dat94] Proceedings, the 10th International Conference on Data Engineering: February 14–18, 1994, Houston, TX, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, February 1994. ISBN 0-8186-5402-3 (case), 0-8186-5400-7 (paperback), 0-8186-5401-5 (microfiche). LCCN QA76.9.D3 I5582 1994. IEEE Computer Society Press order number 5400-02. IEEE Catalog Number 94CH3383-7.

Doerr:1996:UDL

[DL96] H. M. Doerr and B. R. Land. Using a dataflow language and the World Wide Web for scientific visualization. Computer Applications in Engineering Education, 4 (2):161–168, ???? 1996. CODEN CAPEED. ISSN 1061-3773.

Eichinger:1995:PIL

[ECSK95] W. Eichinger, N. Collins, R. Sefecka, and Howard Kraye. Program integrates lidar data for display. Laser Focus World, 31(10): S15–S17, October 1995. CODEN LFWOE8. ISSN 1043-8092.

Earnshaw:1993:ASV

[EW93] R. A. Earnshaw and D. Watson, editors. Animation and scientific visualization: tools and applications. Academic Press, New York, NY, USA, 1993. ISBN 0-12-227745-7. LCCN T385 .A68 1993. Proceedings of the British Computer Society Conference on Animation and Scientific Visualization, December 1992.

Foster:1994:REK

[F^{+}94] M. Foster et al. RCWP expert, a knowledge and GIS based system for selection, evaluation, and design of water quality control practices in agricultural watersheds. In Proceedings GRASS User’s Conference, March 1994, page ?? ????, ????, 1994.

Finsterwalder:1996:AFI

[FG96] R. Finsterwalder and G. Grubel. The ANDECS framework for integrating dynamics modeling, simulation and CACSD tools in interactive loops. In IEE [IEE96], pages 3/1–3/2.

Ford:1994:MSS

[FRN94] R. Ford, S. Running, and R. Nemani. A modular system for scalable ecological modeling. In IEEE C325I Sciences & Engineering, 1, Fall, 1994, pages 32– 44. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994.

Franklin:1994:UEU

[FSW94] S. D. Franklin, A. R. Stubberud, and L. P. Wiedeman, editors. University education uses of visualization in scientific computing: proceedings of the IFIP WG 3.2 Working Conference on Visualization in Scientific Computing, Uses in University Education, Irvine, CA, USA, number A-48 in IFIP Transactions. A, Computer Science and Technology. North-Holland Publishing Co., Amsterdam, The Netherlands, July 1994. ISBN 0-444-81543-0. ISSN 0926-5473. LCCN T385 .I194 1993.

Ford:1995:DIB

[FTD95] R. Ford, D. Thompson, and T. Duce. Data interchange between GIS and visualization environments. In Anonymous [Ano95b], pages 1011–1015 (vol.

2). ISBN 1-882610-15-6. Two volumes.

Garcia:1992:PVS

[GPT92] Armando Garcia, Abraham Peled, and Lloyd A. Treinish. The power visualization system, 1992. 1 videocassette (50 min.).

Gillilan:1995:VEE

[GR95] R. E. Gillilan and D. R. Ripoll.

Visualizing enzyme electrostatics visualization Data Explorer. In Bowie and Olson [BO95], chapter 3, page ?? ISBN 0-201-58157-4. LCCN QD461 .D22 1995. Contents: Introduction — Application visualization system — Quantum chemistry visualization using the AVS Chemistry Viewer — Visualizing enzyme electrostatics with IBM Visualization Data Explorer — Interactive visualization of molecular scenarios with MOLCAD/SYBYL — Molecular surface comparison using Anaconda — Visualization approaches in quantum chemistry using UniChem — Visualizing electrostatic interactions with Insight II — Cerius2 molecular modeling software for industrial applications — The CAChe WorkSystem and molecular mechanics: coordination of DBT on MoS2 — Visualization and representation in chemical education using HyperChem — Visualization supplier reference.

Gillilan:1995:VVR

[GW95] Richard E. Gillilan and Frank Wood. Visualization, virtual reality, and animation with the data flow model of computing. Computer Graphics, 29 (2):55–58, May 1995. CODEN CGRADI, CPGPBZ. ISSN 0097-8930 (print), 1558-4569 (electronic).

Haber:1991:DMS

[H^{+}91] R. Haber et al. A data model for scientific visualization with provisions for regular and irregular grids. In Nielson and Rosenblum [NR91], pages 298–305. ISBN 0-8186-2245-8 (paper), 0-8186-2247-4 (case), 0-8186-2246-6 (microfiche), 0-8186-2245-8 (paperback). LCCN T385 .I192 1991. IEEE Computer Society Press order number 2245. IEEE catalog number 91CH3046-0.

Halatsis:1994:PPA

[Hal94] C. Halatsis, editor. PARLE ’94: parallel architectures and languages Europe: 6th International PARLE Conference, Athens, Greece, July 4–8, 1994: proceedings, volume 817 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CODEN LNCSD9. ISBN 3-540-58184-7 (Berlin), 0-387-58184-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I564 1994.

Haus:1993:VRS

[Hau93] J. Haus. Visualization of real and simulated data in physical oceanography. In Earnshaw and Watson [EW93], pages 225–236.

ISBN 0-12-227745-7. LCCN T385 .A68 1993. Proceedings of the British Computer Society Conference on Animation and Scientific Visualization, December 1992.

Heasley:1994:SCW

[Hea94] Cynthia Heasley, editor. Science and A Changing World: Proceedings of AAAS ’94. (The 160th National Meeting of American Association for the Advancement of Science), San Francisco, February 18–23, 1994, number 94-04S in AAAS Publications. American Association for the Advancement of Science, Washington, DC, USA, February 1994. ISBN ???? LCCN ????

Hay:1993:VTH

[HK93] L. Hay and L. Knapp. Visualization techniques for hydrologic modeling. In Burton [Bur93], pages 3–1–3–8. ISBN ???? LCCN GB701 .W39 no. 93-4018. PB Reports PB95-186615.

Hackstadt:1994:NGP

[HM94] S. T. Hackstadt and A. D. Malony. Next-generation parallel performance visualization: A prototyping environment for visualization development. In Halatsis [Hal94], pages 192–201. CODEN LNCSD9. ISBN 3-540-58184-7 (Berlin), 0-387-58184-7 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I564 1994. URL http:/

/www.cs.uoregon.edu/~hacks/
research/publications.html.

Also, University or Oregon, Deopt. of Computer and Information Science, Technical Report CIS-TR-23, October 1993.

Hackstadt:1995:CSA

[HM95a] S. T. Hackstadt and A. D. Malony. Case study: applying scientific visualization to parallel performance visualization. Proceedings of the SPIE — The International Society for Optical Engineering, 2410:238–247, 1995. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Hackstadt:1995:VBV

[HM95b] Steven T. Hackstadt and Allen D. Malony. Visualization blackboard: Visualizing parallel programs and performance. IEEE Computer Graphics and Applications, 15(4):12–14, July 1995. CODEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Hackstadt:1994:SPV

[HMM94] S. T. Hackstadt, A. D. Malony, and B. Mohr. Scalable performance visualization for dataparallel programs. In SHPCC ’94 [SHP94], pages 342–349. ISBN 0-8186-5680-8 (paper), 0-8186-5681-6 (microfiche). LCCN QA76.5 .S244 1994. URL http:/

                                        /www.cs.uoregon.edu/~hacks/
                                        research/publications.html.

Also, University of Oregon, Dept of Computer and Information Science, Technical Report CIS-TR-94-09, March, 1994.

Heath:1995:VDP

[HMR95] Michael T. Heath, Allen D. Malony, and Diane T. Rover. The visual display of parallel performance data. Computer, 28(11): 21–28, November 1995.

Hostetter:1994:GCS

[Hos94] Carl F. Hostetter, editor. 1994 Goddard Conference on Space Applications of Artificial Intelligence: proceedings of a conference held at NASA Goddard Space Flight Center, Greenbelt, Maryland, May 10–12, 1993 [i.e. 1994], number 3268 in NASA Conference Publication. National Aeronautics and Space Administration, Washington, DC, USA, May 1994. ISSN 0191-7811. LCCN TL 787 G62 1994.

Haines:2000:VCB

[HPM^{+}00] K. G. Haines, B. A. Pearlmutier, J. A. Moya, K. Edlund, and T. P. Caudell. Visualizing communication between neurons in the lamina ganglionaris of Musca domestica. In S.-I. Amari, C. L. Giles, M. Gori, and V. Piuri, editors, International Joint Conference on Neural Networks, 2000. IJCNN 2000, Proceedings of the IEEE-INNS-ENNS, 24–27 July 2000, volume 3, pages 53– 58. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7695-0619-4. LCCN QA76.5 .I558 2000; QA76.87.I57

IBM:1993:IVDa

[IBM93a] IBM Corporation, Yorktown Heights, NY, USA. IBM visualization Data Explorer, third edition, February 1993. various pp.

IBM:1993:IVDb

[IBM93b] IBM Corporation, Yorktown Heights, NY, USA. IBM visualization Data Explorer, fourth edition, October 1993. various pp.

IEE:1996:ICA

[IEE96] IEE, editor. IEE Colloquium on Advances in Computer-Aided Control System Design (Digest), volume 96. IEE, London, UK, 1996.

Kalvin:1992:VAR

[K^{+}92] A. Kalvin et al. Visualization in anthropology: Reconstruction of human fossils from multiple pieces. In Kaufman and Nielson [KN92a], pages 404–409. ISBN 0-8186-2897-9 (case), 0-8186-2896-0 (microfiche), 0-8186-2897-9. LCCN T385 .I192 1992. IEEE Computer Society Press order number 2897. IEEE catalog no. 92CH3201-1.

King:1995:ATV

[KCWG95] B. King, P. Collins, E. Wolanski, and D. Galloway. Animation techniques for visualizing coastal flow dynamics. In Spaulding and Cheng [SC95], pages 192– 203. ISBN 0-7844-0165-9. LCCN GC96.5 .E772 1996.

Kalvin:1995:RHF

[KDH95] Alan D. Kalvin, David Dean, and Jean-Jacques Hublin. Reconstruction of human fossils. IEEE Computer Graphics and Applications, 15(1):12–15, January 1995. CO-DEN ICGADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Kalvin:1994:VRR

[KDHB94] Alan Kalvin, David Dean, Jean-Jacques Hublin, and Marc Braun. Virtual reality reconstruction of early homo sapiens fossils from morocco. In Heasley [Hea94], pages 152–?? ISBN ???? LCCN ???? Abstract number 355.

King:1996:FVR

[KG96] E. King and Richard Gillilan. The frontiers of virtual reality and visualization in biochemical research. Scientific Computing & Automation, 8:20–22, July 1996. CODEN SCOAEG. ISSN 0891-9003.

Kelly:1993:JMF

[KL93] Gary G. Kelly and Anthony J.

Lewis, editors. Joint meeting — February 1993, New Orleans, LA, number 1 in ACSM ASPRS Annual Convention. American Congress on Surveying and Mapping. American Society for Photogrammetry and Remote Sensing, Falls Church, VA, USA, February 1993. ISBN 0-944426-93-X (set), 0-944426-94-8 (vol. 1), 0-944426-95-6 (vol. 2), 0-944426-96-4 (vol. 3). LCCN ???? Three volumes. Vol. 1: ACSM technical papers / Gary G. Kelly, editor; Vol. 2: ASPRS technical papers, remote sensing / Anthony J. Lewis, editor; Vol. 3: AS-PRS technical papers, GIS, photogrammetry and modeling / Anthony J. Lewis, editor.

Kaufman:1992:VPO

[KN92a] Arie E. Kaufman and Gregory M. Nielson, editors. Visualization ’92: proceedings, October 19–23, 1992 Boston, MA, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2897-9 (case), 0-8186-2896-0 (microfiche), 0-8186-2897-9. LCCN T385 .I192 1992. IEEE Computer Society Press order number 2897. IEEE catalog no. 92CH3201-1.

Kaufman:1992:SVR

[KN92b] Arie E. Kaufman and Gregory M. Nielson, editors. Visualization ’92: proceedings, October 19– 23, 1992, Boston, Massachusetts. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2897-9. LCCN T385 .I192 1992. IEEE catalog number 92CH3201-1.

Knapp:1993:TAG

[Kna93] L. Knapp. Task analysis and geographic visualization. In Kelly and Lewis [KL93], pages 190– 199. ISBN 0-944426-93-X (set), 0-944426-94-8 (vol. 1), 0-944426-95-6 (vol. 2), 0-944426-96-4 (vol. 3). LCCN ???? Three volumes. Vol. 1: ACSM technical papers / Gary G. Kelly, editor; Vol. 2: ASPRS technical papers, remote sensing / Anthony J. Lewis, editor; Vol. 3: ASPRS technical papers, GIS, photogrammetry and modeling / Anthony J. Lewis, editor.

Knapp:1994:TAA

[Kna94] L. Knapp. A task analysis approach to visualization of geographic data. In Nyerges et al. [N^{+}94], pages 355–372. ISBN 0-7923-3595-3. LCCN G70.2.C63 1995.

Lucus:1992:ASV

[L^{+}92] B. Lucus et al. An architecture for a scientific visualization system. In Kaufman and Nielson [KN92a], pages 107–113. ISBN 0-8186-2897-9 (case), 0-8186-2896-0 (microfiche), 0-8186-2897-9. LCCN T385 .I192 1992. IEEE Computer Society Press order number 2897. IEEE catalog no. 92CH3201-1.

Lucas:1992:ASV

[LAC^{+}92] B. Lucas, G. D. Abram, N. S.

Collins, D. A. Epstein, D. L. Gresh, and K. P. McAuliffe. An architecture for a scientific visualization system. In Kaufman and Nielson [KN92b], pages 107–114. ISBN 0-8186-2897-9. LCCN T385 .I192 1992. IEEE catalog number 92CH3201-1.

Land:1994:TCG

[Lan94] B. R. Land. Teaching computer graphics and scientific visualization using the dataflow, block diagram language Data Explorer. In Franklin et al. [FSW94], pages 33– 36. ISBN 0-444-81543-0. ISSN 0926-5473. LCCN T385 .I194 1993.

Liddell:1996:HCN

[LCHS96] Heather Mary Liddell, A. Colbrook, B. Hertzberger, and P. Sloot, editors. Highperformance computing and networking: international conference and exhibition, HPCN EUROPE 1966, Brussels, Belgium, April 15–19, 1996: proceedings, volume 1067 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61142-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .H52 1996.

Land:1993:SVC

[LG93] B. R. Land and R. E. Gillilan. Scientific visualization of chemical systems. In Supercomputing ’93 [Sup93], pages 296–300. ISBN 0-8186-4342-0, 0-8186-4340-4, 0-8186-4341-2 (microfiche). ISSN 1063-9535. LCCN QA76.5 .S894 1993. IEEE Computer Society Press order number 4340-02. IEEE catalog number 93CH3342-3. ACM order number 415932.

Larkin:1996:VLS

[LGH96] S. Larkin, A. J. Grant, and W. T. Hewitt. “Vipar” libraries to support distribution and processing of visualization datasets. In Liddell et al. [LCHS96], pages 711– 721. ISBN 3-540-61142-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .H52 1996.

Lee:1995:MRE

[LNH^{+}95] K. Lee, Jun Ni, T. Halverson, E. van Wyk, and J. R. Brown. Is an MVE the right environment for your visualization application? Computer Graphics, 29(2):49–51, May 1995. CO-DEN CGRADI. ISSN 0097-8930 (print), 1558-4569 (electronic).

Lucas:1992:SVR

[Luc92] B. Lucas. A scientific visualization renderer. In Kaufman and Nielson [KN92a], pages 227–234. ISBN 0-8186-2897-9 (case), 0-8186-2896-0 (microfiche), 0-8186-2897-9. LCCN T385 .I192 1992. IEEE Computer Society Press order number 2897. IEEE catalog no. 92CH3201-1.

Ma:1998:SVD

[MJS98] H. Ma, K. W. Jones, and E. A.

Stern. Scientific visualization and data modeling of scattered sediment contaminant data in New York/New Jersey estuaries. In D. Ebert, H. Rushmeier, and H. Hagen, editors, Proceedings: Visualization ’98, 18–23 October 1998, pages 67–470, 569. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-9176-X. LCCN T385 .I42 1998. IEEE catalog number 98CB36276.

Nyerges:1994:CAH

[N^{+}94] Timothy L. Nyerges et al., editors. Cognitive aspects of human-computer interaction for geographic information systems:

NATO Advanced research workshop — March 20–25, 1994, Palma de Mallorca, Spain, number 83 in NATO ASI Series D Behavioural and Social Sciences. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, March 1994. ISBN 0-7923-3595-3. LCCN G70.2.C63 1995.

Nielson:1993:VAC

[NB93] Gregory M. Nielson and Dan Bergeron, editors. Visualization ’93: proceedings, October 25–29, 1993, San Jose, CA, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 1993. ISBN 0-8186-3942-3 (case), 0-8186-3940-7 (paper), 0-8186-3941-5 (microfiche). ISSN 1070-2385. LCCN T385 .I192 1993.

Nielson:1991:VPO

[NR91] Gregory M. Nielson and Lawrence J. Rosenblum, editors. Visualization ’91: proceedings, October 22– 25, 1991, San Diego, CA, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 1991. ISBN 0-8186-2245-8 (paper), 0-8186-2247-4 (case), 0-8186-2246-6 (microfiche), 0-8186-2245-8 (paperback). LCCN T385 .I192 1991. IEEE Computer Society Press order number 2245. IEEE catalog number 91CH3046-0.

Nielson:1995:VPO

[NS95] Gregory M. Nielson and Deborah E. Silver, editors. Visualization ’95 proceedings, October 29–November 3, 1995, Atlanta, GA, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 1995. ISBN 0-7803-3127-3, 0-8186-7187-4, 0-7803-3128-1. ISSN 1070-2385. LCCN T385 .I198 1995. IEEE Computer Society Press order number PR07187. IEEE Catalog Number 95CB35835.

OConnor:1994:VAC

[OS94] Lisa O’Connor and Penny Storms, editors. Visualization ’94: proceedings, October 17–21, 1994, Washington, DC, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 1994. ISBN 0-8186-6627-7 (case), 0-8186-6626-9 (microfiche). ISSN 1070-2385. LCCN T385 .I192 1994. Fifth Annual IEEE Conference on Visualization. IEEE catalog number 94CH35707.

PhillipsMahoney:1996:ASM

[Phi96] Diana Phillips Mahoney. The art and science of medical visualization — in conjunction with dramatic advances in scanning technologies, computer-aided visualization tools are. Computer Graphics World, 19(7):25– ??, July 1996. CODEN CG-WODH. ISSN 0271-4159.

Pickover:1994:APG

[Pic94] C. A. Pickover. Automatic parallel generation of Aeolian fractals on the IBM Power Visualization System. Computers and Graphics, 18(3):407–416, May– June 1994. CODEN COGRD2. ISSN 0097-8493 (print), 1873-7684 (electronic).

Paddeu:1995:CMG

[PMA^{+}95] G. Paddeu, D. H. Mosley, J.- M. Andre, J. G. Fripiat, and D. P. Vercauteren. Computeraided molecular graphics for the interpretation of electronic structure calculations on polymers. In Scateni [Sca95], pages 79– 91. ISBN 981-02-2461-3. LCCN Q174.S337 1995.

Ribarsky:1992:ODV

[R^{+}92] W. Ribarsky et al. Objectoriented, dataflow visualization systems — A paradigm shift? In Kaufman and Nielson [KN92a], pages 384–387. ISBN 0-8186-2897-9 (case), 0-8186-2896-0 (microfiche), 0-8186-2897-9. LCCN T385 .I192 1992. IEEE Computer Society Press order number 2897. IEEE catalog no. 92CH3201-1.

Rogowitz:1994:HVV

[RA94] Bernice E. Rogowitz and Jan P. Allebach, editors. Human vision, visual processing, and digital display V: 8–10 February 1994, San Jose, California, number 2179 in Proceedings — SPIE, The International Society for Optical Engineering. Society of Photooptical Instrumentation Engineers (SPIE), Bellingham, WA, USA, February 1994. ISBN 0-8194-1474-3. LCCN ???? Proceedings of a conference held February, 1994, in San Jose, CA, USA.

Ribarsky:1992:POD

[RBM^{+}92] W. Ribarsky, B. Brown, T. Myerson, S. Smith, and L. Treinish. Panel on object-oriented, dataflow visualization systems – A paradigm shift? In Kaufman and Nielson [KN92a], pages 384– 387. ISBN 0-8186-2897-9 (case), 0-8186-2896-0 (microfiche), 0-8186-2897-9. LCCN T385 .I192 1992. IEEE Computer Society Press order number 2897. IEEE catalog no. 92CH3201-1.

Rhyne:1992:CSV

[RBP^{+}92] T.-M. Rhyne, M. Botts, E. Pepke, W. Hibbard, and L. A. Treinish. Case study on visualization requirements in the atmospheric and environmental sciences. In Kaufman and Nielson [KN92a], page ?? ISBN 0-8186-2897-9 (case), 0-8186-2896-0 (microfiche), 0-8186-2897-9. LCCN T385 .I192 1992. IEEE Computer Society Press order number 2897. IEEE catalog no. 92CH3201-1.

Rogowitz:1993:ARVb

[RT93a] B. E. Rogowitz and L. A. Treinish.

An architecture for rule-based visualization. In Nielson and Bergeron [NB93], pages 236–243 (or 236–244??). ISBN 0-8186-3942-3 (case), 0-8186-3940-7 (paper), 0-8186-3941-5 (microfiche). ISSN 1070-2385. LCCN T385 .I192

Rogowitz:1993:APR

[RT93b] Bernice E. Rogowitz and Lloyd A. Treinish. An architecture for perceptual rule-based visualization. In Nielson and Bergeron [NB93], pages 236–243. ISBN 0-8186-3942-3 (case), 0-8186-3940-7 (paper), 0-8186-3941-5 (microfiche). ISSN 1070-2385. LCCN T385 .I192 1993.

Rogowitz:1993:DSP

[RT93c] Bernice E. Rogowitz and Lloyd A. Treinish. Data structures and perceptual structures. In Allebach and Rogowitz [AR93], pages 600– 612. ISBN 0-8194-1146-9. ISSN 0361-0748. LCCN ????

Rogowitz:1993:ARVa

[RT93d] Bernice Ellen Rogowitz and Lloyd A. Treinish. An architecture for rule-based visualization. Research report RC 19196 (83625), IBM T. J. Watson Research Center, Yorktown Heights, NY, USA, October 4, 1993. 14 pp.

Rogowitz:1994:UPR

[RT94] Bernice E. Rogowitz and Lloyd A. Treinish. Using perceptual rules in interactive visualization. In Rogowitz and Allebach [RA94], pages 287–297. ISBN 0-8194-1474-3. LCCN ???? Proceedings of a conference held February, 1994, in San Jose, CA, USA.

Rogowitz:1996:PVH

[RT96] Bernice E. Rogowitz and Lloyd A. Treinish. PRACTICAL VISU-ALIZATION: How not to lie with visualization. Computers in Physics, 10(3):268–??, May 1996. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Rogowitz:1998:DVE

[RT98] B. E. Rogowitz and L. A. Treinish.

Data visualization: the end of the rainbow. IEEE Spectrum, 35(12): 52–59, December 1998. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

Safran:1994:SAS

[Saf94] Charles Safran, editor. Seventeenth Annual Symposium on Computer Applications in Medical Care. Patient-Centered Computing. McGraw-Hill, New York, NY, USA, 1994. ISBN 0-07-001502-3. LCCN W3 SY5074K 1993.

Stiles:1994:ACW

[SB94] Joel Stiles and Thomas Bartol, Jr.

About the cover: Walking randomly in a synapse. IEEE Computational Science & Engineering, 1(3):4–5, Fall 1994. CODEN IS-CEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Spaulding:1995:ECM

[SC95] Malcolm L. Spaulding and Ralph T. Cheng, editors. Estuarine and coastal modeling: proceedings of the 4th international conference: October 26–28, 1995, San Diego, California. American Society of Civil Engineers, New York, NY, USA, 1995. ISBN 0-7844-0165-9. LCCN GC96.5 .E772 1996.

Scateni:1995:SVP

[Sca95] Riccardo Scateni, editor. Scientific Visualization ’95: proceedings of the International Symposium, Cagliari, Italy, 27–29 September, 1995. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1995. ISBN 981-02-2461-3. LCCN Q174.S337 1995.

Schenker:1993:SFV

[Sch93] Paul S. Schenker, editor. Sensor fusion VI: 7–8 September 1993, Boston, MA, USA, number 2059 in Proceedings SPIE: The International Society For Optical Engineering. Society of Photooptical Instrumentation Engineers (SPIE), Bellingham, WA, USA, September 1993. ISBN 0-8194-1324-0. ISSN 0361-0748. LCCN TS510.S63 v.2059.

IEEE:1994:PSH

[SHP94] Proceedings of the Scalable High-Performance Computing Conference, May 23–25, 1994, Knoxville, TN, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, May 1994. ISBN 0-8186-5680-8 (paper), 0-8186-5681-6 (microfiche). LCCN QA76.5 .S244 1994. IEEE catalog number 94TH0637-9.

Song:1995:HVT

[SN95] Deyang Song and M. Norman. 4d2 — a high-performance visualization tool for CFD and particle data visualization. Proceedings of the SPIE — The International Society for Optical Engineering, 2410:139–146, 1995. CO-DEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

IEEE:1993:SCN

[Sup93] Proceedings, Supercomputing ’93:

Portland, Oregon, November 15– 19, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4342-0, 0-8186-4340-4, 0-8186-4341-2 (microfiche). ISSN 1063-9535. LCCN QA76.5 .S894 1993. IEEE Computer Society Press order number 4340-02. IEEE catalog number 93CH3342-3. ACM order number 415932.

Treinish:1996:VBW

[TC96] L. A. Treinish and Z. D. Christidis. Visualization blackboard: Weather forecasting for the 1996 Olympics. IEEE Computer Graphics and Applications, 16(4): 10–13, July 1996. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Treinish:1991:CVT

[TG91] L. A. Treinish and C. Goettsche.

Correlative visualization techniques for multidimensional data. IBM Journal of Research and Development, 35(1/2):184–204, January/March 1991. CODEN IB-MJAE. ISSN 0018-8646 (print), 2151-8556 (electronic).

Treinish:1993:DMS

[TKFG93] L. Treinish, R. Kulkarni, M. Folk, and G. Goucher. Data models, structures, and access software for scientific visualization. In Nielson and Bergeron [NB93], pages 355–361. ISBN 0-8186-3942-3 (case), 0-8186-3940-7 (paper), 0-8186-3941-5 (microfiche). ISSN 1070-2385. LCCN T385 .I192 1993.

Treinish:1992:CDA

[Tre92a] L. A. Treinish. Correlative data analysis in the earth sciences. IEEE Computer Graphics and Applications, 13(3):10– 12, May 1992. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Treinish:1992:PIV

[Tre92b] Lloyd A. Treinish. Climatology of global stratospheric ozone (1979 through 1991). In Proceedings IEEE Visualization ’92, video supplement, page ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, October 1992. ISBN ???? LCCN ????

Treinish:1992:DMM

[Tre92c] Lloyd A. Treinish. Data management methods for scientific visualization. Research report RC 18315 (80219), IBM T. J. Watson Research Center, Yorktown Heights, NY, USA, September 9, 1992. 26 pp.

Treinish:1993:UPD

[Tre93a] L. A. Treinish. Unifying principles of data management for scientific visualization. In Earnshaw and Watson [EW93], pages 141–169.

ISBN 0-12-227745-7. LCCN T385 .A68 1993. Proceedings of the British Computer Society Conference on Animation and Scientific Visualization, December 1992.

Treinish:1993:VSO

[Tre93b] L. A. Treinish. Visualization of stratospheric ozone depletion and the polar vortex. In Nielson and Bergeron [NB93], pages 391–396. ISBN 0-8186-3942-3 (case), 0-8186-3940-7 (paper), 0-8186-3941-5 (microfiche). ISSN 1070-2385. LCCN T385 .I192 1993.

Treinish:1993:VTC

[Tre93c] L. A. Treinish. Visualization techniques for correlative data analysis in the earth and space sciences. In Earnshaw and Watson [EW93], pages 193–204. ISBN 0-12-227745-7. LCCN T385 .A68 1993. Proceedings of the British Computer Society Conference on Animation and Scientific Visualization, December 1992.

Treinish:1993:BLD

[Tre93d] Lloyd A. Treinish. Browsing large data archives: Visualization for data management. Ibm research internal report, IBM Corporation, Yorktown Heights, NY, USA, March 1993. ???? pp.

Treinish:1993:CGS

[Tre93e] Lloyd A. Treinish. Climatology of global stratospheric ozone (1979 through 1991). ACM SIGGRAPH Video Review, 93:??, August 1993.

Treinish:1993:CVE

[Tre93f] Lloyd A. Treinish. Correlative visualization in the earth and space sciences. In Earnshaw and Watson [EW93], pages 193–204. ISBN 0-12-227745-7. LCCN T385 .A68 1993. Proceedings of the British Computer Society Conference on Animation and Scientific Visualization, December 1992.

Treinish:1993:IMD

[Tre93g] Lloyd A. Treinish. Inside multidimensional data. Byte Magazine, 18(4):??, April 1993. CO-DEN BYTEDJ. ISSN 0360-5280 (print), 1082-7838 (electronic).

Treinish:1993:VAE

[Tre93h] Lloyd A. Treinish. Visualization and analysis of environmental data. In Visualization ’93 Course #2: The Process of Visualizing Environmental Data Sets, page ?? ????, ????, October 1993. ISBN ???? LCCN ????

Treinish:1993:VAG

[Tre93i] Lloyd A. Treinish. Visualization and analysis of global data. In ACM SIGGRAPH ’93 Course #71: Visualizing Planet Earth, page ?? ????, ????, August 1993. ISBN ???? LCCN ????

Treinish:1994:IASb

[Tre94a] L. A. Treinish. Interactive archives of scientific data. In Hostetter [Hos94], pages 3–16. ISSN 0191-7811. LCCN TL 787 G62 1994.

Treinish:1994:IASc

[Tre94b] L. A. Treinish. Interactive archives of scientific data. Telematics and Informatics, 11(4): 333–??, November 1994. CO-DEN TEINEG. ISSN 0736-5853 (print), 1879-324X (electronic).

Treinish:1994:SRE

[Tre94c] L. A. Treinish. Severe rainfall events in northwestern peru: Visualization of scattered meteorological data. In O’Connor and Storms [OS94], pages 350– 354. ISBN 0-8186-6627-7 (case), 0-8186-6626-9 (microfiche). ISSN 1070-2385. LCCN T385 .I192 1994. Fifth Annual IEEE Conference on Visualization. IEEE catalog number 94CH35707.

Treinish:1994:VSMb

[Tre94d] L. A. Treinish. Severe rainfall events in northwestern peru: Visualization of scattered meteorological data. In O’Connor and Storms [OS94], pages 350– 354. ISBN 0-8186-6627-7 (case), 0-8186-6626-9 (microfiche). ISSN 1070-2385. LCCN T385 .I192 1994. Fifth Annual IEEE Conference on Visualization. IEEE catalog number 94CH35707.

Treinish:1994:IASa

[Tre94e] Lloyd A. Treinish. Interactive archives of scientific data. Research report RC 19539 (84992), IBM T. J. Watson Research Center, Yorktown Heights, NY, USA, April 28, 1994. 13 pp.

Treinish:1994:STD

[Tre94f] Lloyd A. Treinish. Solutions and techniques for data management, visual display and examination of large scientific data sets. global ozone distribution. In ACM SIG-GRAPH ’94 Course #27: Visualizing and Examining Large Scientific Data Sets – A Focus on the Physical and Natural Sciences, page ?? ACM Press, New York, NY 10036, USA, 1994. ISBN ???? LCCN ????

Treinish:1994:VSMa

[Tre94g] Lloyd A. Treinish. Visualization of scattered meteorological data: study of severe rainfall events in northwestern peru. Research report RC 19537 (84985), IBM T. J. Watson Research Center, Yorktown Heights, NY, USA, March 1994. 6 + 2 pp.

Treinish:1994:VSO

[Tre94h] Lloyd A. Treinish. Visualization of stratospheric ozone and atmospheric dynamics (1991). ACM SIGGRAPH Video Review, 105: ??, July 1994.

Treinish:1994:VID

[Tre94i] Lloyd A. Treinish. Visualizations illuminate disparate data sets in the earth sciences. Computers in Physics, 8(6):664–??, November/December 1994. CO-DEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

Treinish:1995:GAA

[Tre95a] L. A. Treinish. A generalized approach to analysis and visualization of atmospheric, earth and environmental sciences data. In Anonymous [Ano95a], pages 19– 24.

Treinish:1995:IAE

[Tre95b] L. A. Treinish. Interactive archives of earth science data. In Anonymous [Ano95a], pages 13– 18.

Treinish:1995:VSM

[Tre95c] Lloyd A. Treinish. Visualization of scattered meteorological data. IEEE Computer Graphics and Applications, 15(4):20– 26, July 1995. CODEN IC-GADZ. ISSN 0272-1716 (print), 1558-1756 (electronic).

Watson:1994:IE

[Wat94] D. Watson. Intrepid Explorer. In Computer Graphics Development International, page ?? Comhill Publications Limited, London, UK, 1994. ISBN 1-85938-013-1. LCCN ????

Watson:19xx:ANE

[Watxx] D. Watson. Adding non-defined element types to IBM Data Explorer. Report 279, IBM UK Scientific Center, IBM Hursley, Winchester, UK, 19xx. ???? pp.